Two opposite angles of a parallelogram are (4x-5) and (50-x).Find the measure of each angle of the parallelogram.
Answer by jsmallt9(3759)

The opposite angles in a parallelogram are congruent. So:
(4x-5) = (50-x)

Now we solve for x. Adding x:
5x-5 = 50
Add 5:
5x = 55
Divide by 5:
x = 11

From x we can find the two angles:
4(11)-5 = 44-5 = 39
The two given angles are 39 degrees.

The other two opposite angles will be congruent, too. And they will be supplementary to the first two angles. If we call these other two angles y then:
29 + y = 180
Subtracting 29 from each side:
y = 151
So the other two angles are 151 degrees each.